[ ]/45 = 16/40 = 24/[ ]
 [ ]/45 = 16/40 = 24/[ ]
 [ ]/45 = 16/40 = 24/[ ]

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Consider the first two ratios [ ]/45 = 16/40 Let us assume the missing number be P Then, P/45 = 16/40
By cross multiplication we get,

P = (16 × 45)/40
P = 720/40
P = 72/4
Divide both numerator and denominator by 4. P = 18

Therefore, [18]/45 = 16/40
Now, consider the last two ratios, 16/40 = 24/[ ] Let us assume the missing number be Q,
Then, 16/40 = 24/Q
By cross multiplication we get,
Q = (24 × 40)/16
Q = 960/16
Divide both numerator and denominator by 16. Q = 60
Therefore, 16/40 = 24/[60]


Thanks